OpenStudy (anonymous):

Sam did 480 J of work when moving a bag with a mass of 38 kg. He did the work in 4 minutes. What was the power output? 442 W 20 W 120 W 2 W

OpenStudy (anonymous):

Power = work/time

OpenStudy (s3a):

ΔP = ΔW/Δt where the change of work is in joules and the change in time is in seconds. You can define the initial work and time to be 0 so that you don't need to think in terms of change but just of how things are in the end. That is, in the specific interval of time in question, 480 J of energy was expended in 4 minutes * 60 seconds/1minute = 240 seconds. Therefore, ΔP = ΔW/Δt = (480 J)/(240 s) = 2 W is your answer. I don't think you know Calculus yet but, for future reference, you can replace Δ by d such that ΔP = ΔW/Δt becomes dP = dW/dt; the difference is that d is an infinitely small change which is useful for getting instantaneous results whereas the Δ can be any amount of change. For example, thinking in terms of instantaneous behaviour, in other words, the smallest possible change, hence the d, of speed/velocity, (which is not what your problem is about) about how fast someone is driving a car can mean "right now," this person is going this fast whereas the average change, Δ, means that, on average, in this person's entire journey, this person was going that fast. Hope this helps. P.S. I'm very sleepy as I write this so forgive any mistakes.